OpenStudy (nuccioreggie):

Which expression is equivalent to 54 ⋅ 5–6? 5–10 56 fraction 1 over 5 to the power 2 fraction 1 over 5 to the power negative10

OpenStudy (nuccioreggie):

@whpalmer4

Vocaloid (vocaloid):

rule of exponents: a^b/a^c = a^(b-c)

OpenStudy (nuccioreggie):

@Triciaal

OpenStudy (nuccioreggie):

@welshfella

OpenStudy (nuccioreggie):

Please help dont leave

OpenStudy (whpalmer4):

is that \[54\cdot 5^{-6}\]

OpenStudy (whpalmer4):

or \[5^4\cdot 5^{-6}\]?

OpenStudy (whpalmer4):

I'm guessing the latter

OpenStudy (welshfella):

I'm a bit confused by the question is it